### v3

# SpikeNAS: A Fast Memory-Aware Neural Architecture Search Framework for Spiking Neural Network-based Autonomous Agents

Abstract
Autonomous mobile agents (e.g., UAVs and UGVs) are typically expected to incur low power/energy consumption for solving machine learning tasks (such as object recognition), as these mobile agents are usually powered by portable batteries. These requirements can be fulfilled by Spiking Neural Networks (SNNs), since their bio-inspired spike-based operations offer high accuracy and ultra low-power/energy computation. Currently, most of the SNN architectures are derived from Artificial Neural Networks whose neurons’ architectures and operations are different from SNNs, or developed without considering memory budgets from the underlying processing hardware of autonomous mobile agents. These limitations hinder SNNs from reaching their full potential in accuracy and efficiency. Toward this, we propose SpikeNAS, a novel fast memory-aware neural architecture search (NAS) framework for SNNs that quickly finds an appropriate SNN architecture with high accuracy under the given memory budgets from autonomous mobile agents. To do this, our SpikeNAS employs several key steps: analyzing the impacts of network operations on the accuracy, enhancing the network architecture to improve the learning quality, and developing a fast memory-aware search algorithm. The experimental results show that our SpikeNAS improves the searching time and maintains high accuracy as compared to state-of-the-art while meeting the given memory budgets (e.g., 4.4x faster search with 1.3% accuracy improvement for CIFAR100, using an Nvidia RTX 6000 Ada GPU machine), thereby quickly providing the appropriate SNN architecture for the memory-constrained autonomous mobile agents.
I Introduction
In recent years, autonomous mobile agents (e.g., UAVs, UGVs, and mobile robots) are expected to incorporate neural network (NN) algorithms [1, 2, 3], as these algorithms achieve state-of-art accuracy when solving diverse machine learning (ML) tasks, such as image classification [4], object recognition [5], and continual learning [6]. These mobile agents are also expected to incur low power/energy consumption when executing NN algorithms, because they are usually powered by portable batteries [2][3]. These requirements can be fulfilled by Spiking Neural Networks (SNNs), as their bio-inspired spike-based operations offer high accuracy while incurring ultra low-power/energy computation [7]. Therefore, many research works are carried out to maximize the SNN potentials, ranging from SNN architecture developments to its deployments for diverse ML-based applications [8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. In fact, different ML-based applications usually have different accuracy requirements, which often necessitate different memory costs to facilitate the corresponding SNN architectures. For instance, a bigger SNN is typically developed to achieve a higher accuracy, but at the cost of a larger memory footprint, as shown in Fig. 1(a). Furthermore, a larger memory footprint leads to a higher energy consumption, as the memory accesses still dominate the overall processing energy of SNNs; see Fig. 1(b). Hence, the development of SNN architectures that can achieve the expected accuracy with acceptable memory footprint is needed to support the widespread use of SNNs for autonomous agents.
Figure 1: (a) Autonomous mobile agents are typically expected to achieve high accuracy with low power and energy-efficient processing to complete ML-based tasks. (b) Accuracy and memory of different SNN models for CIFAR10: CIFARNet1 [18], CIFARNet2 [19], ResNet11 [20], ResNet19 [21], AutoSNN [13]; adapted from studies in [13]. (c) Breakdown of energy consumption for SNN processing by different hardware platforms: PEASE [22], TrueNorth [23], and SNNAP [24]; adapted from studies in [25].
Currently, most of the SNN architectures are derived from Artificial Neural Networks (ANNs), whose neurons’ architectures and operations are different from SNNs [13]. SNNs accumulate and generate temporal information in form of spikes using Leaky Integrate-and-Fire (LIF) neurons, while ANNs perform multiplication-and-accumulation (MAC) and activation using Rectified Linear Unit (ReLU). Such ANN-based SNN architectures may lose unique features from SNN computation models (e.g., temporal information between spikes), thereby leading to sub-optimal accuracy [14]. Moreover, the SNN architectures are still developed without considering memory budgets from the underlying processing hardware, which is important to ensure the applicability of the SNN architecture for the given hardware platform. For instance, most of the SNN models require more than 20M of parameters to achieve 93% accuracy for CIFAR10, as shown in Fig. 1(a), which is memory consuming and inefficient for resource-constrained autonomous mobile agents. These limitations hinder the SNNs from reaching their full potential in accuracy and efficiency for autonomous mobile agents. However, manually designing the desired SNN architecture is a laborious task and time consuming, thus requiring an alternative solution that quickly provides an SNN architecture that is suitable for the given accuracy and memory constraints.
Research Problem: How to automatically and quickly find an SNN architecture that achieves acceptable accuracy under the given memory constraints for autonomous mobile agents. An efficient solution to this problem will ease SNN developments for autonomous mobile agents, thereby enabling the widespread use of SNNs for diverse ML-based tasks [26].
I-A State-of-the-Art and Their Limitations
To address the research problem, previous works employ neural architecture search (NAS) to automatically find SNN architectures that can achieve high accuracy. Currently, NAS for SNNs is still at the early stage, hence only a few studies found in the literature [13][14][27, 28, 29], and the prominent ones are discussed in the following.
  • AutoSNN [13]: It aims at finding an SNN architecture with high accuracy and low number of spikes by employing a NAS with training approach. First, it trains a super-network that includes all of the possible candidates of SNN architectures, then performs a search to find a subset of network architecture that achieves the target. However, this NAS approach requires a huge searching time as well as a huge memory and energy consumption, as it needs to perform training for a large network before searching and determining the desired SNN architecture.
  • SNASNet [14]: It aims at finding an SNN architecture that achieves high accuracy by employing a NAS without training approach to reduce the searching and training time. It is performed by exploring the architecture candidates of the cell (i.e., a directed acyclic graph that represents network topology and operations; see Fig. 2a), evaluating the score of the investigated architecture, then training the architecture with the highest score [14]. However, it has several limitations as follows.
    • It employs a random search mechanism to determine the cell architecture, hence the effectiveness and duration of the search depend on the number of search iteration.
    • It has a fixed number of cells (i.e., 2 cells), and it applies the same architecture for all cells, thereby reducing the possibility to find a better SNN architecture (e.g., higher accuracy).
Moreover, these state-of-the-art works also do not consider memory constraints from the target hardware in their search algorithms, thereby limiting the efficiency of the generated SNN architectures when deployed on the target hardware. To identify and better understand the limitations of state-of-the-art (i.e., a NAS without training approach), we perform an experimental case study and discuss it in Section I-B.
I-B Case Study and Research Challenges
Figure 2: (a) A cell is a directed acyclic graph whose edge corresponds to an operation from the pre-defined operations [30]. (b) Results of accuracy across different numbers of search iteration and different operation types.
We study the impacts of search iterations and operation types on the accuracy considering a NAS without training approach using the search space and mechanism of SNASNet, i.e, exploring the topology and operations of a cell architecture in Fig. 2(a) using a random search. To evaluate the impact of search iterations, the state-of-the-art work employs 5000x iterations and the investigated one employs 4096x. To evaluate the impact of operation types, the state-of-the-art work employs 5 operation types in Fig. 2(a), while the investigated one employs 4 operation types (i.e., without Zeroize). Details of the experimental setup are provided in Section IV. The experimental results are presented in Fig. 2(b), from which we draw the following key observations.
  • A smaller number of search iterations may lead to comparable accuracy across different training epochs, meaning that there is a possibility to maintain accuracy with a smaller search space.
  • A smaller number of operation types may lead to comparable (or better) accuracy across different training epochs, meaning that there is a possibility to maintain (or improve) accuracy with a smaller search space.
These observations expose several research challenges and requirements (outlined below) in devising solutions for the targeted research problem.
  • The solution should quickly perform the search to minimize the searching time.
  • The solution should be able to guide the search mechanism towards finding an SNN architecture that maintains high accuracy, while avoiding redundancy.
  • The solution should incorporate the memory constraints into its searching algorithm to ensure the applicability and efficiency of the SNN deployments on the targeted hardware platforms.
I-C Our Novel Contributions
Towards this, we propose SpikeNAS, a novel NAS framework that can quickly find the appropriate SNN architecture with high accuracy under the given memory budgets from autonomous mobile agents. Our SpikeNAS employs several key steps as the following (see an overview in Fig. 3, and the details in Fig. 4).
  • Analysis of the Impacts of Network Operations: It aims to understand the significance of each pre-defined operation type. It is done by removing the investigated operation type from the search space and then observing the output accuracy.
  • Network Architecture Enhancements: It aims to maintain (or improve) accuracy of the network architecture, while reducing the searching time. It is done by optimizing the cell operation types and the number of cells in the network architectures.
  • A Fast Memory-aware Search Algorithm: It aims to find an SNN architecture that achieves high accuracy and meets the memory budget, by performing search for each individual cell, while monitoring the memory cost of the investigated architecture.
Key Results: We evaluate our SpikeNAS framework using a Python-based implementation on an Nvidia RTX 6000 Ada GPU machine. The experimental results show that our SpikeNAS improves the searching time and maintains high accuracy as compared to state-of-the-art while meeting the memory budgets, e.g., SpikeNAS improves the searching time by 4.4x and accuracy by 1.3% under 2M parameters for CIFAR100.
Figure 3: Overview of our novel contributions (in blue boxes).
Figure 4: Overview of our SpikeNAS framework, and its key steps for generating an SNN architecture that offers high accuracy while meeting the given memory constraint from an autonomous mobile agent. The novel contributions are highlighted in blue boxes.
II Preliminaries
II-A Spiking Neural Networks (SNNs)
SNNs are considered the third generation of NN computation model [31]. SNNs typically employ a LIF neuron model [32, 33, 34, 35, 36], whose neuronal dynamics can be formulated as follows.
$\begin{split}\tau_{leak}\frac{dV_{mem}(t)}{dt}&=-(V_{mem}(t)-V_{reset})+x(t)\end{split}$ (1)
$\begin{split}\text{if}\;\;\;V_{mem}&\geq V_{th}\;\;\;\text{then}\;\;\;V_{mem}\leftarrow V_{reset}\end{split}$ (2)
$V_{mem}$ denotes the neurons’ membrane potential at timestep-$t$, $V_{reset}$ denotes the neurons’ reset potential, and $\tau_{leak}$ denotes the time constant of membrane potential decay. Meanwhile, $x$ denotes the inputs, which can be stated in a discrete time form as the following.
$\begin{split}x^{l}[t]=\sum_{i}wgh_{i}^{l}\;\delta_{i}^{l-1}[t]+b^{l}\end{split}$ (3)
$x$ denotes the neuron input for layer-$l$ at timestep-$t$, $b$ denotes the bias, and $wgh$ denotes the weight at index-$i$. Meanwhile, $\delta$ denotes the input spike from previous layer ($l-1$) at index-$i$ and timestep-$t$.
II-B NAS without Training
NAS without training approach was proposed in ANN domain [37], and then leveraged for SNN domain [14]. Its key idea is to quickly measure the capability of a given network to reach high accuracy at early stage, thereby avoiding the training phase to evaluate the network. It is motivated by the observation that an architecture with distinctive representations across different samples is likely to achieve high accuracy after training [37]. To measure the representation capability of a network, the state-of-the-art [14] employs the following key steps (our SpikeNAS framework also follows the same steps for evaluating the network).
  • It employs LIF neuron to divide the input space into active and inactive regions. If $V_{mem}$ reaches $V_{th}$ then the neuron is mapped to 1, otherwise 0. Then, the activity of LIF neurons in each layer is represented as binary vector $f$.
  • It feeds $S$ samples in a mini-batch to the network. Then, a kernel matrix $\textbf{K}_{H}$ for each layer is constructed by computing the Hamming distance $H(f_{i},f_{j})$ between different samples $i$ and $j$; see Eq. 4. Here, $F$ is the number of LIF neurons in the given layer, while $\alpha$ is the factor to normalize high sparsity from LIF neurons.
  • It calculates the score of the architecture candidate using Eq. 5. Then, the candidate with the highest score is selected for training.
$\textbf{K}_{H}=\begin{pmatrix}F-\alpha H(f_{1},f_{1})&\cdots&F-\alpha H(f_{1},f_{S})\\ \vdots&\ddots&\vdots\\ F-\alpha H(f_{S},f_{1})&\cdots&F-\alpha H(f_{S},f_{S})\end{pmatrix}$ (4)
$\begin{split}score=\log(\det\;\lvert\sum_{l}\textbf{K}_{H}^{l}\rvert)\end{split}$ (5)
II-C Neural Cell-based Strategy for NAS
Neural cell (or simply cell)-based strategy has been widely used for NAS in ANNs. The motivation of cell-based NAS is to provide a unified benchmark for NAS algorithms [30]. Its approach is to search for the connection topology and the corresponding operations inside each cell; see Fig. 2(a). A cell typically consists of 4 nodes and 5 pre-defined operation types: No-Connection (Zeroize), Skip-Connection (SkipCon), 1x1 Convolution (1x1Conv), 3x3 Convolution (3x3Conv), and 3x3 Average Pooling (3x3AvgPool). The connection between 2 nodes employs a specific pre-defined operation type. Therefore, there are 15625 architecture candidates in each cell. The state-of-the-art (i.e., SNASNet) employs a 2-cell SNN architecture for its NAS technique; see its macro-architecture in Fig. 5. In this work, we also employ the same macro-architecture for our SpikeNAS framework, and focus on the feedforward topology as it is widely used in the SNN community.
III SpikeNAS Framework
To address the above-discussed research problem and challenges, we propose a novel SpikeNAS framework; see the detailed flow in Fig. 4. Its key steps are discussed in the Section III-A - Section III-C, consecutively.
III-A Analysis of the Impact of Operation Types
Discussion in Section I-B suggests that employing all pre-defined operations might not lead to optimal solution. Therefore, to understand the significance of each pre-defined operation, we perform an experimental case study. Here, we remove a specific operation type, then perform NAS considering a macro-architecture in Fig. 5 with 5000x search iteration. Since we employ the NAS without training approach, then this step is performed before the search process and the training phase. Experimental results are shown in Fig. 6, from which we draw the following key observations.
  • 1
    Removing the 3x3Conv operation from the cells significantly reduces the accuracy, as this operation employs 3x3 kernel size for extracting important features from input feature maps. Therefore, removing the 3x3Conv make the network difficult to perform learning properly. Therefore, the 3x3Conv operation should be kept in the search space.
  • 2
    Removing the Zeroize, 1x1Conv, 3x3AvgPool operations from the cells do not change the accuracy significantly. Therefore, these operations may be removed from the search space.
  • 3
    Removing the SkipCon operation from the search space also notably reduces the accuracy for CIFAR100, as this operation provides feature maps from previous layer to ensure the preservation of important information. This observation also aligns with the observation from studies in [38]. Therefore, the SkipCon operation should be kept in the search space.
  • 4
    In general, the order of significance of the operations from the highest to the lowest: (1) 3x3Conv, (2) SkipCon, and (3) Zeroize, 1x1Conv, and 3x3AvgPool. We consider Zeroize, 1x1Conv, and 3x3AvgPool in the same significance level as they have comparable impact on the accuracy. These operations may be removed to reduce the search space based on their significance.
Figure 5: The macro-architecture for a cell-based NAS in the state-of-the-art (i.e., employing 2 cells).
Figure 6: Results of removing different types of operations from the search space of a cell architecture. These results are for CIFAR100, and results for CIFAR10 show similar trends.
Afterward, these observations are then leveraged as guidance for enhancing the NAS, so that it can quickly find and generate SNN architecture that has significant operations or characteristics to achieve high accuracy, while considering the given memory constraints.
III-B Network Architecture Enhancements
III-B1 Optimization of the Operation Types in the Neural Cells
In Section III-A, we observe that some operations can be removed from the search space to improve the searching time without sacrificing accuracy. Hence, we optimize the number of operation types based on their significance on the accuracy. We propose two variants of operation sets that can be considered in the search space to provide multiple optimization options (see an overview in Fig. 7), as follows.
  • First operation set: SkipCon, 3x3Conv, and 3x3AvgPool. We select 3x3AvgPool to reduce the size of feature maps, while minimizing the compute and memory requirements and ensuring the connectivity to the next layer.
  • Second operation set: SkipCon and 3x3Conv.
Figure 7: The pre-defined cell operation types in SpikeNAS. First operation set consists of 3 operation types, while second operation set consists of 2 operation types.
III-B2 SNN Macro-Architecture Adjustments
To increase the applicability and efficiency of SNN deployments for diverse ML-based applications, we propose adjustments for the SNN macro-architecture. Apart from the 2-cell SNN architecture, we also propose different variants of network architectures based on the number of cells (i.e., 1 cell and 3 cells), as discussed in the following.
  • A 1-Cell Architecture: It aims at reducing the memory footprint, but might reduce the accuracy as compared to the 2-cell architecture; see an illustration in Fig. 8. Its key idea is removing one cell from the 2-cell architecture; see an overview in Fig. 9(a).
  • A 3-Cell Architecture: It aims at improving the accuracy, but might incur a larger memory footprint as compared to the 2-cell architecture; see an illustration in Fig. 8. Its key idea is adding one cell to the 2-cell architecture; see an overview in Fig. 9(b).
In this manner, our SpikeNAS framework provides multiple design options to trade-off accuracy, memory footprint, and searching time.
Figure 8: The accuracy vs memory trade-off among the 1-cell, 2-cell, and 3-cell architectures.
Figure 9: The proposed variations of the network architectures in SpikeNAS with (a) 1 cell and (b) 3 cells.
III-C A Fast Memory-aware Search Algorithm
To synergistically perform architecture search while considering the memory constraint, we propose a fast memory-aware search algorithm. The key ideas of our search algorithm are discussed in the following, while the corresponding pseudo-code is shown in Alg. 1.
  • Each cell has an individual search to determine its architecture, thereby exploring more candidates from different combinations of architectures across multiple cells (Alg. 1: line 2).
  • Investigation of redundant architectures for each cell is minimized to improve the effectiveness of our architecture search. To do this, we propose an in-cell architecture search (Alg. 1: line 3-14). Variables $con_{xy}$ with $xy$$\in\{01,02,03,12,13,23\}$ in Alg. 1: line 9-14 correspond to the type of operation that connect node-$X$ to node-$Y$; see Fig. 7. Then, the cell architecture is constructed once $con_{xy}$ conections are defined (Alg. 1: line 15).
  • All cells are initiated with the same architecture (Alg. 1: line 16-17). Then, the architecture candidates are explored for each cell, starting from the second cell to the last one, consecutively (Alg. 1: line 18-19). Hence, the search space is reduced while keeping good architectures from previous cells. Then, the network architecture is constructed once all cells are defined (Alg. 1: line 21).
  • Information of the memory constraint is included into the search process (Alg. 1: line 22). It is used to select which architectures that should be evaluated, hence filtering out the ones that are not suitable and optimizing the searching time even further (Alg. 1: line 23-27). To enable this, an analytical model of memory footprint is developed and employed into the search. Details of the proposed analytical model is discussed in the following.
Algorithm 1 Our Fast Memory-aware Search Algorithm
0:  (1) Pre-defined number of neural cells ($C$); (2) Neural cells: object names ($cell_{1}$, …, $cell_{C}$), neural cell architecture ($arch$) ;(3) Memory constraint or budget ($mem_{budget}$);(4) Pre-defined number of operation types ($T$); (5) Operation codes: SkipCon (0), 3x3Conv (1), and 3x3AvgPool (2);
0:  Generated SNN architecture ($net\_arch$); BEGIN Initialization:
1:  $score_{h}$=-1000; Process:
2:  for ($i$ = 1; $i$¡($C$+1); $i$++) do
3:     for ($c01$=0; $c01$¡$T$; $c01$++) do
4:        for ($c02$=0; $c02$¡$T$; $c02$++) do
5:           for ($c03$=0; $c03$¡$T$; $c03$++) do
6:              for ($c12$=0; $c12$¡$T$; $c12$++) do
7:                 for ($c13$=0; $c13$¡$T$; $c13$++) do
8:                    for ($c23$=0; $c23$¡$T$; $c23$++) do
9:                       $con_{01}=c01$;
10:                       $con_{02}=c02$;
11:                       $con_{03}=c03$;
12:                       $con_{12}=c12$;
13:                       $con_{13}=c13$;
14:                       $con_{23}=c23$;
[…]
16:                       if ($i$==1) then
17:                          $cell_{1},...,cell_{N}\leftarrow arch$;
18:                       else
19:                          $cell_{i}\leftarrow arch$;
20:                       end if
21:                       $net$ = create$(cell_{1},...,cell_{C})$;
22:                       $mem_{cost}$= memcalc($net$);
23:                       if ($mem_{cost}\leq mem_{budget}$then
24:                          $score=$ eval$(net)$;
25:                          if ($score>score_{h}$then
26:                             $score_{h}=score$;
27:                             $net\_arch=net$;
28:                          end if
29:                       end if
30:                    end for
31:                 end for
32:              end for
33:           end for
34:        end for
35:     end for
36:  end for
37:  return  $net\_arch$; END
Analytical Model for Memory Footprint: We focus on the number of parameters in the SNN architecture ($N_{param}$) to represent the memory footprint, which can be calculated using Eq. 6-7.
$\begin{split}N_{param}=\sum_{l}(N_{wgh}^{l}+N_{b}^{l})\end{split}\vspace{-0.3cm}$ (6)
$\begin{split}N_{wgh}=h_{wgh}\cdot w_{wgh}\cdot c_{wgh}\cdot r_{wgh}\end{split}$ (7)
$N_{wgh}^{l}$ and $N_{b}^{l}$ denotes the number of weights and bias in layer-$l$, respectively. The 2-dimensional size of filters in the given layer is represented with $h_{wgh}$ for height and $w_{wgh}$ for width. Meanwhile, $c_{wgh}$ and $r_{wgh}$ denote the number of channel and the number of filters in the given layer, respectively. Furthermore, the memory footprint can also be represented in bit or byte format if required ($M_{param}$), by considering the bit precision ($bit$) as stated in Eq. 8.
$\begin{split}M_{param}=N_{param}\cdot bit\end{split}$ (8)
IV Evaluation Methodology
[…]
For evaluation and validation of the proposed concepts, we develop a Python-based implementation using the SpikingJelly library [40] on a multi-GPU machine (i.e., an Nvidia RTX 6000 Ada GPU machine with 48GB GDDR6 memory), and consider CIFAR10 and CIFAR100 as the workloads. We perform up to 300 training epochs with surrogate gradient learning, while evaluating the test accuracy gradually to see the learning curve of the investigated network architecture. These experiments provide information of accuracy, searching time, and memory footprint.
Figure 10: The experimental setup used for the evaluation.
TABLE I: Different scenarios for SpikeNAS with 3 operation types: SkipCon, 3x3Conv, and 3x3AvgPool.
Number of Cells
1 2 3
SpikeNAS_1C3O
SpikeNAS_1C3O_M
SpikeNAS_2C3O
SpikeNAS_2C3O_M
SpikeNAS_3C3O
SpikeNAS_3C3O_M
TABLE II: Different scenarios for SpikeNAS with 2 operation types: SkipCon and 3x3Conv.
Number of Cells
1 2 3
SpikeNAS_1C2O
SpikeNAS_1C2O_M
SpikeNAS_2C2O
SpikeNAS_2C2O_M
SpikeNAS_3C2O
SpikeNAS_3C2O_M
V Results and Discussion
Figure 11: Results of the accuracy for (a) CIFAR10 and (b) CIFAR100.
Figure 12: Results of the memory footprint normalized to the state-of-the-art work SNASNet_2C5O for (a) CIFAR10, and (b) CIFAR100.
Figure 13: Results of the searching time normalized to the state-of-the-art work SNASNet_2C5O for (a) CIFAR10, and (b) CIFAR100.
V-A Maintaining the High Accuracy
Fig. 11(a) and Fig. 11(b) present the experimental results for the accuracy for CIFAR10 and CIFAR100, respectively.
The 2-Cell Architectures: SNASNet achieves 92.58% of accuracy for CIFAR10 and 69.18% for CIFAR100. Our SpikeNAS_2C3O achieves 93.06% of accuracy for CIFAR10 and 71.62% for CIFAR100, while SpikeNAS_2C2O achieves 93.4% of accuracy for CIFAR10 and 71.26% for CIFAR100; see 1 and 2. These results indicate that, our SpikeNAS improves the accuracy as compared to the state-of-the-art, since SpikeNAS only considers operations with high significance to the accuracy. Moreover, all of these accuracy improvements can be achieved by networks from SpikeNAS with less than 10M parameters. When the memory constraints are given, our SpikeNAS_2C3O_M achieves 90.96% of accuracy for CIFAR10 and 70.52% for CIFAR100, while SpikeNAS_2C2O_M achieves 92.12% of accuracy for CIFAR10 and 68.02% for CIFAR100; see 1 and 2. These results show that, our SpikeNAS maintains (or improves in some cases) the accuracy as compared to the state-of-the-art, while ensuring that the generated SNN architecture meets the memory budget. The reason is that, SpikeNAS employs operations with high significance to the accuracy and incorporates the memory budget in its search process. Note, memory evaluation is discussed in Sec. V-B.
The 1-Cell and 3-Cell Architectures (Variants): Our SpikeNAS_1C3O achieves 91.29% of accuracy for CIFAR10 and 64.79% for CIFAR100, while SpikeNAS_1C2O achieves 91.60% of accuracy for CIFAR10 and 65.25% for CIFAR100; see 3 and 4. Furthermore, our SpikeNAS_3C3O achieves 93.51% of accuracy for CIFAR10 and 73.07% for CIFAR100, while SpikeNAS_3C2O achieves 93.43% of accuracy for CIFAR10 and 73.25% for CIFAR100; see 3 and 4. These results indicate that, in general, the 1-cell architectures offer a lower accuracy as compared to the 2-cell architectures, as the 1-cell architecture have fewer operations to extract meaningful features from the inputs. Meanwhile, 3-cell architectures can offer a higher accuracy as compared to the 2-cell architectures, as the 3-cell architecture have more operations to extract meaningful features from the inputs. When the memory constraints are given, SpikeNAS_1C3O_M and SpikeNAS_1C2O_M have the similar results to their unconstrained counterparts (i.e., SpikeNAS_1C3O and SpikeNAS_1C2O) as the memory footprints of the unconstrained ones are already within the given memory budgets. Meanwhile, SpikeNAS_3C3O_M and SpikeNAS_3C2O_M have lower accuracy than their unconstrained counterparts (i.e., SpikeNAS_3C3O and SpikeNAS_3C2O), since the memory constraint adds a limitation for finding a combination of cell architectures that offers high accuracy.
These results highlight that, our SpikeNAS can generate SNN architectures that maintain or improve the accuracy as compared to the state-of-the-art.
V-B Ensuring Memory-aware SNN Generation
Fig. 12(a) and Fig. 12(b) present the experimental results for the memory footprint for CIFAR10 and CIFAR100, respectively. For CIFAR10, SNASNet incurs 6% memory overhead than the given budget, while SpikeNAS can save the memory footprint up to 43% from the state-of-the-art; see 5. If we consider high accuracy, then we can select SpikeNAS_2C2O_M with 92.12% accuracy and 30% memory saving from the state-of-the-art (24% memory saving from the given budget); see 6. Meanwhile, for CIFAR100, SNASNet incurs 23% memory overhead than the given budget, while SpikeNAS can save the memory footprint up to 72% from the state-of-the-art; see 7. If we consider high accuracy, then we can select SpikeNAS_2C3O_M with 70.52% accuracy and 38% memory saving from the state-of-the-art (15% memory saving from the given budget); see 8. These results highlight that, SpikeNAS can generate SNN architectures that meet the given memory constraints due to its memory-aware search algorithm, enabling efficient deployments of SNN models on the target hardware platforms.
V-C Reducing the Searching Time
Fig. 13(a) and Fig. 13(b) present the experimental results for the searching time for CIFAR10 and CIFAR100, respectively. In both cases of workloads (CIFAR10 and CIFAR100), our SpikeNAS offers significant improvements for the searching time as compared to the state-of-the-art. The reason is that, SpikeNAS employs fewer operation types that have significant impact on the accuracy, performs a consecutive search across the existing cells, and incorporates a memory constraint into the search process, thereby leading to a smaller search space and a faster searching time than the state-of-the-art. If we consider high accuracy, acceptable memory cost, and fast searching time, then we can select SpikeNAS_2C2O_M for CIFAR10 with 92.12% accuracy, 23% memory saving from the given memory budget, and 347.7x faster search than the state-of-the-art; see 9. For CIFAR100, we can select SpikeNAS_2C3O_M with 70.52% accuracy, 14% memory saving from the given memory budget, and 4.4x faster search than the state-of-the-art; see 10. These results highlight that, our SpikeNAS can quickly generate SNN architectures that offer high accuracy and meet the memory constraints, hence enabling efficient design automation for developing SNN-based autonomous mobile agents.
VI Conclusion
We propose the SpikeNAS framework to quickly find an SNN architecture that provides high accuracy under a memory constraint from an autonomous mobile agent. SpikeNAS employs an analysis of the impact of operations on the accuracy, architecture enhancements, and a NAS that is guided by memory budget and consecutive searches across cells. The experimental results show that, SpikeNAS maintains the accuracy and improves the searching time as compared to the state-of-the-art with a smaller memory cost. For instance, our SpikeNAS improves the searching time by 4.4x and accuracy by 1.3% under 2M parameters for CIFAR100. In this manner, our SpikeNAS framework enables the efficient SNN deployments for autonomous mobile agents.
Acknowledgment
This work has been supported in parts by the NYUAD Center for Cyber Security (CCS), funded by Tamkeen under the NYUAD Research Institute Award G1104, and Center for Artificial Intelligence and Robotics (CAIR), funded by Tamkeen under the NYUAD Research Institute Award CG010.